OpenStudy (anonymous):

Write the equation of the line that passes through (-6,2) with slope 2/3.

OpenStudy (anonymous):

use \[y-y_1=m(x-x_1)\] with \[x_1=-6,y_1=2,m=\frac{2}{3}\]

OpenStudy (anonymous):

that is the method. if you just want the answer, cheat http://www.wolframalpha.com/input/?i=line+that+passes+through+%28-6%2C2%29+with+slope+2%2F3